OpenStudy (anonymous):

Plz Help ASAP gold medal will be awarded.for the function f(x)=x+5, what is the ordered pair for the point on the graph when x=4w

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

solving f(x) will give you a y value so it is the same as saying y = x+5 in the function you are asked to evaluate y at x = 4w inset 4w in place of x in function. Proper way to ask is f(4w) = x + 5 That means you are evaluating for y at the value of 4w f(4w) = (4w)+5 y = 4w +5 at x=4w

OpenStudy (anonymous):

oh (4w, 4w+5) is the answer choice that i can find

OpenStudy (anonymous):

plz help